OpenStudy (anonymous):

Is the exact value of cot 930 square root of 3/3

OpenStudy (anonymous):



OpenStudy (anonymous):

It's actually \[\cot {930} ???\] in radians?

OpenStudy (anonymous):

No 930 is in degrees

OpenStudy (anonymous):

Oh, okay, that makes it simpler, trigonometric functions are circular functions, so, every 360º they repeat themselves, that means that you'll have to calculate cot 210, but cot (x) is equal to cos(x)/sen(x), now, cos(210º)=-cos(30º) and sen(210º)=-sin (30º) Now, \[\cos(30º)=\dfrac {\sqrt{3}} {2} \] and \[\sin(30º)=0.5\] Do your maths c: